OpenStudy (anonymous):

What is the exact volume of a sphere with a diameter of 18 cm?

OpenStudy (welshfella):

V = (4/3) pi r^3 where r = radius ( radius = 1/2 diameter)

OpenStudy (anonymous):

what?